OpenStudy (anonymous):

Solve for h. a=ch+f 1. h=a+f/c 2. h=a-f/c 3. h=a/c -f 4. h=a-c/f

OpenStudy (alexandervonhumboldt2):

subtract f from both sides: a-f=ch+f-f a-f=ch divide both sides by c: ch/c=(a-f)/c h=(a-f)/c

OpenStudy (anonymous):

thanks

OpenStudy (alexandervonhumboldt2):

np